# Adversarial Training on Purification (AToP): Advancing Both Robustness and Generalization

### 10 Mar 2024 v2

@@ 4 Experiments
In this section, we conduct extensive experiments on CIFAR-10, CIFAR-100 and ImageNette across various transforms, classifier models and purifier models on attack benchmarks. Our findings demonstrate that the proposed method achieves state-of-the-art results and exhibits generalization ability against unseen attacks. Specifically, compared to against AutoAttack $l_{\infty}$ on CIFAR-10, our method improves the robust accuracy by 13.28% on GAN-based purifier model and 14.45% on AE-based purifier model, respectively. Furthermore, we empirically study the random transforms and disccuss their impacts on robust accuracy. Finally, we provide visual examples (more details can be found in supplementary materials) to provide a more comprehensive understanding of our results.
@@ 4.1 Experimental setup
Adversarial attacks: We evaluate our method against various attacks: We utilize AutoAttack $l_{\infty}$ and $l_{2}$ threat models (Croce & Hein, 2020) as one benchmark, which is a powerful attack that combines both white-box and black-box attacks, and utilizes Expectation Over Time (EOT) (Athalye et al., 2018b) to tackle the stochasticity introduced by random transforms. To consider unseen attacks without $l_{p}$-norm, we utilize spatially transformed adversarial examples (StAdv) (Xiao et al., 2018) for validation. Additionally, we generate adversarial examples using some standard methods including Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015), Projected Gradient Descent (PGD) (Madry et al., 2018b) and Carlini-Wagner (CW) Attack (Carlini & Wagner, 2017).
@@
Training details: After experimental testing, we have determined the hyperparameters: gaussian standard $\sigma=0.25$, mask size $p=P/8$ ($P$ is side length of input $x$), mask number $N=4$, missing rate $r=0.25$, and weight $\lambda=0.1$. All experiments presented in the paper are conducted under these hyperparameters and done using the NVIDIA RTX A5000 in PyTorch v1.13.1 (Paszke et al., 2019).
@@ 4.2 Comparison with state-of-the-art methods
We evaluate our method against AutoAttack $l_{\infty}$ and $l_{2}$ threat models and compare its robustness to state-of-the-art methods listed in RobustBench (Croce et al., 2021). All experiments presented in this section are utilizing $\rm{RT_{2}}$ and GAN-based model, which yielded the best results in the experiment.
@@
Result analysis on CIFAR-100: Table 4 shows the performance of the defense methods against AutoAttack $l_{\infty}$ ($\epsilon=8/255$) threat model on CIFAR-100 with WideResNet-28-10 and WideResNet-70-16 classifier models. Our method outperforms all other methods without extra data in terms of both standard accuracy and robust accuracy. Specifically, compared to the second-best method, our method improves the robust accuracy by 4.37% on WideResNet-28-10 and by 1.77% on WideResNet-70-16, respectively. Table 5 shows the performance of the models against AutoAttack $l_{\infty}$ ($\epsilon=8/255$) and the observations are basically consistent with CIFAR-10, demonstrating the effectiveness of our method in helping robust classification across different datasets.
Table 6: Standard accuracy and robust accuracy against AutoAttack $l_{\infty}$ ($\epsilon=8/255$), $l_{2}$ ($\epsilon=1$) and StAdv non-$l_{p}$ ($\epsilon=0.05$) threat models on CIFAR-10 with ResNet-50 classifier model. We utilize GAN-based model with $\rm{RT_{2}}$ and all settings follow the same as Laidlaw et al. (2021).
@@
4.3 Defense against unseen attacks
As previously mentioned, mainstream adversarial training (AT) methods have shown limitations in defending against unseen attacks (Laidlaw et al., 2021; Dolatabadi et al., 2022), even can only defense against known attacks. To demonstrate the generalization ability of our method with $\rm{RT_{2}}$ and GAN, we conduct experiments in which defense methods against attacks with varying constraints (including AutoAttack $l_{\infty}$, $l_{2}$ and StAdv non-$l_{p}$ threat models) on CIFAR-10 with ResNet-50 classifier model.
Table 6 shows that AT method can achieve robustness against particular attacks seen during training (as indicated by Acc. with underscore) but struggles to generalize robustness to unseen attacks. The intuitive idea is to train with AT across all attacks to obtain a robust model. However, it is challenging to train such a model due to the inherent differences among attacks. Furthermore, the emergence of new attack techniques continually makes it impractical to train the model to defense against all attacks. We combine AT and AP as a new defense method, which outperforms all other methods in terms of both standard accuracy and robust accuracy. Specifically, compared to the second-best method, our method improves the robust accuracy by 1.2%, 2.5% and 1.4% on AutoAttack $l_{\infty}$, $l_{2}$ and StAdv non-$l_{p}$, respectively. The results demonstrate that our method achieves state-of-the-art results and exhibits generalization ability against unseen attacks.
@@ 4.4 Ablation studies
We conduct ablation studies to validate the effectiveness of factors within our method defense against of some standard attacks. Considering the robustness misestimation caused by obfuscated gradients of the purifier model, we use BPDA (Athalye et al., 2018a) following the setting by Yang et al. (2019), which approximating the gradient of purifier model as 1 during the backward pass. Additionally, we use EOT Athalye et al. (2018b) following the setting by Lee & Kim (2023) with 20 iterations.
@@
Figure 3: The purified images obtained by clean (Top) and adversarial examples (Bottom) with different random transforms.
Table 8: Standard accuracy and robust accuracy against PGD+EOT $l_{2}$ threat ($\epsilon=1.0$) on CIFAR-10 with ResNet-18 classifier model. AToP Standard PGD-10 PGD-20 $\rm{RT_{1}}$ 93.36 20.31 15.04 $\times$ 93.36 27.34 22.46 $\rm{RT_{2}}$ 84.18 67.19 63.09 $\times$ 90.04 69.92 68.75 $\rm{RT_{3}}$ 75.98 28.52 26.17 $\times$ 80.02 46.48 41.60
Figure 3: The purified images obtained by clean (Top) and adversarial examples (Bottom) with different random transforms.
@@
For the more comprehensive evaluation, we follow the guidance of Lee & Kim (2023) to evaluate our method against PGD+EOT with $l_{2}$ threat ($\epsilon=1.0$). Table 8 shows the performance of the defense model with different transforms against PGD-10 and PGD-20. The results show that our method significantly outperforms the method without AToP in terms of both standard accuracy and robust accuracy, demonstrating the effectiveness of our method in enhancing the performance of the purifier model in robust classification.
@@
Figure 4: Standard accuracy and robust accuracy of the GAN-based purifier model trained with clean examples and adversarial examples, respectively, on (a) CIFAR-10 and (b) ImageNette. The dashed line represents the accuracy of standard training w/o attacks. (c) Standard accuracy and robust accuracy of our method with adversarial trained ResNet-18 on CIFAR-10.
@@
Thus far, we have conducted extensive experiments across various attacks, classifiers, generator-based purifiers, datasets, and RT settings, which have consistently demonstrated the effectiveness of our method in enhancing robust classification.
Limitations: Given that our method involves the fine-tuning of the generator model, the complexity of the generator model has a direct impact on the computational cost of AToP. Initially, we explore fine-tuning the diffusion model to improve its robustness. However, as purified images need to be fed into the classifier, generating 1000 images and fine-tuning the diffusion model requires 144 minutes. In contrast, with the same setting, the GAN-based purifier model only requires 62 seconds.
@@ 5 Conclusion
In the paper, we propose a novel defense technique called AToP that utilizes adversarial training to optimize the purifier model. To demonstrate the robustness of our method, we conduct extensive experiments on CIFAR-10, CIFAR-100 and ImageNette. In defense of various attacks within multiple transforms, classifier and purifier architectures, our method consistently demonstrated that it achieves state-of-the-art results and exhibits generalization ability against unseen attacks. Ablation studies further underscore that our method significantly improves the performance of the purifier model for robust classification. Despite the large improvements achieved, our method has a major limitation: AToP requires training on the purifier model, and as the complexity of the purifier model increases, so does the training cost. However, exploring this direction presents an intriguing research opportunity for developing a more efficient defense method by combining AT and AP.

### 15 Mar 2024 v3

@@ 4 Experiments
In this section, we conduct extensive experiments on CIFAR-10, CIFAR-100 and ImageNette across various transforms, classifier models and purifier models on attack benchmarks. Our findings demonstrate that the proposed method achieves state-of-the-art results and exhibits generalization ability against unseen attacks. Specifically, compared to against AutoAttack $l_{\infty}$ on CIFAR-10, our method improves the robust accuracy by 13.28% on GAN-based purifier model and 14.45% on AE-based purifier model, respectively. Furthermore, we empirically study the random transforms and discuss their impacts on robust accuracy. Finally, we provide visual examples (more details can be found in supplementary materials) to provide a more comprehensive understanding of our method.
@@ 4.1 Experimental setup
Adversarial attacks: We evaluate our method against various attacks: We utilize AutoAttack $l_{\infty}$ and $l_{2}$ threat models (Croce & Hein, 2020) as one benchmark, which is a powerful attack that combines both white-box and black-box attacks, and utilizes Expectation Over Time (EOT) (Athalye et al., 2018b) to tackle the stochasticity introduced by random transforms. To consider unseen attacks without $l_{p}$-norm, we utilize spatially transformed adversarial examples (StAdv) (Xiao et al., 2018) for validation. Additionally, we generate adversarial examples using some standard methods, including Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015), Projected Gradient Descent (PGD) (Madry et al., 2018b) and Carlini-Wagner (CW) (Carlini & Wagner, 2017) attacks.
@@
Training details: After experimental testing, we have determined the hyperparameters: Gaussian standard deviation $\sigma=0.25$, mask size $p=P/8$ ($P$ is side length of input $x$), mask number $N=4$, missing rate $r=0.25$ and weight $\lambda=0.1$. All experiments presented in the paper are conducted under these hyperparameters and performed by NVIDIA RTX A5000 Graphics Card with 24GB GDDR6 GPU memory, CUDA v11.7 and cuDNN v8.5.0 in PyTorch v1.13.11 (Paszke et al., 2019).
@@ 4.2 Comparison with state-of-the-art methods
We evaluate our method against AutoAttack $l_{\infty}$ and $l_{2}$ threat models and compare its robustness to state-of-the-art methods listed in RobustBench (Croce et al., 2021). All experiments presented in this section utilize $\rm{RT_{2}}$ and GAN-based model, which yielded the best results in the experiment.
@@
Result analysis on CIFAR-100: Table 4 shows the performance of the defense methods against AutoAttack $l_{\infty}$ ($\epsilon=8/255$) threat model on CIFAR-100 with WideResNet-28-10 and WideResNet-70-16 classifier models. Our method outperforms all other methods without extra data in terms of both standard accuracy and robust accuracy. Specifically, compared to the second-best method, our method improves the robust accuracy by 4.37% on WideResNet-28-10 and by 1.77% on WideResNet-70-16, respectively. Table 5 shows the performance of the models against AutoAttack $l_{\infty}$ ($\epsilon=8/255$) and the observations are basically consistent with CIFAR-10, demonstrating the effectiveness of our method in enhancing robust classification across different datasets.
Table 6: Standard accuracy and robust accuracy against AutoAttack $l_{\infty}$ ($\epsilon=8/255$), $l_{2}$ ($\epsilon=1$) and StAdv non-$l_{p}$ ($\epsilon=0.05$) threat models on CIFAR-10 with ResNet-50 classifier model. We utilize GAN-based model with $\rm{RT_{2}}$, and all settings follow the same as Laidlaw et al. (2021).
@@
4.3 Defend against unseen attacks
As previously mentioned, the mainstream adversarial training (AT) methods have shown limitations in defending against unseen attacks (Laidlaw et al., 2021; Dolatabadi et al., 2022), even can only defend against known attacks. To demonstrate the generalization ability of our method with $\rm{RT_{2}}$ and GAN, we conduct experiments in which defense methods are evaluated against attacks with varying constraints (including AutoAttack $l_{\infty}$, $l_{2}$ and StAdv non-$l_{p}$ threat models) on CIFAR-10 with ResNet-50 classifier model.
Table 6 shows that AT method can achieve robustness against particular attacks seen during training (as indicated by Acc. with underscore) but struggles to generalize robustness to unseen attacks. The intuitive idea is to apply AT across all different attacks to obtain a robust model. However, it is challenging to train such a model due to the inherent differences among all sorts of attacks. Furthermore, the emergence of new attack techniques continually makes it impractical to train the model to defend against all attacks. We combine AT and AP as a new defense method, which outperforms all other methods in terms of both standard accuracy and robust accuracy. Specifically, compared to the second-best method, our method improves the robust accuracy by 1.2%, 2.5% and 1.4% on AutoAttack $l_{\infty}$, $l_{2}$ and StAdv non-$l_{p}$, respectively. The results demonstrate that our method achieves state-of-the-art results and exhibits generalization ability against unseen attacks.
@@ 4.4 Ablation studies
We conduct ablation studies to validate the effectiveness of factors within our method in defending against more standard attacks. Considering the robustness misestimation caused by obfuscated gradients of purifier model, we use BPDA (Athalye et al., 2018a) following the setting by Yang et al. (2019), which approximates the gradient of purifier model as 1 during backward pass. Additionally, we use EOT (Athalye et al., 2018b) following the setting by Lee & Kim (2023) with 20 iterations.
@@
Figure 3: The purified images are obtained for clean (Top) and adversarial examples (Bottom) with different random transforms.
Table 8: Standard accuracy and robust accuracy against PGD+EOT $l_{2}$ threat ($\epsilon=1.0$) on CIFAR-10 with ResNet-18 classifier model. AToP Standard PGD-10 PGD-20 $\rm{RT_{1}}$$\times$ 93.36 20.31 15.04 93.36 27.34 22.46 $\rm{RT_{2}}$$\times$ 84.18 67.19 63.09 90.04 69.92 68.75 $\rm{RT_{3}}$$\times$ 75.98 28.52 26.17 80.02 46.48 41.60
Figure 3: The purified images are obtained for clean (Top) and adversarial examples (Bottom) with different random transforms.
@@
For the more comprehensive evaluation, we follow the guidance of Lee & Kim (2023) to evaluate our method against PGD+EOT with $l_{2}$ threat ($\epsilon=1.0$). Table 8 shows the performance of the defense model with different transforms, and our method still significantly outperforms the method without AToP in terms of both standard accuracy and robust accuracy, demonstrating the effectiveness of our method in enhancing the performance of the purifier model in robust classification.
@@
Figure 4: Standard accuracy and robust accuracy of the GAN-based purifier model trained with clean examples and adversarial examples, respectively, on (a) CIFAR-10 and (b) ImageNette. The dashed line represents the accuracy of standard training w/o attacks. (c) Standard accuracy and robust accuracy of our method with adversarially trained ResNet-18 on CIFAR-10.
@@
So far, we have conducted extensive experiments across various attacks, classifiers, generator-based purifiers, datasets and RT settings, which have consistently demonstrated the effectiveness of our method in acquiring the robust purifier model for enhancing robust classification.
Limitations: Given that our method involves the fine-tuning of the generative model, the complexity of the generative model has a direct impact on the computational cost of AToP. Initially, we explore fine-tuning the diffusion model to improve its robustness. However, as purified images need to be fed into the classifier, generating 1000 images and fine-tuning the diffusion model requires 144 minutes. In contrast, with the same setting, the GAN-based purifier model only requires 62 seconds.
@@ 5 Conclusion
In the paper, we propose a novel defense technique called AToP that utilizes adversarial training to acquire the robust purifier model. To demonstrate the robustness of our method, we conduct extensive experiments on CIFAR-10, CIFAR-100 and ImageNette. In defense of various attacks under multiple transforms, classifier and purifier architectures, our method consistently achieves state-of-the-art results and exhibits generalization ability against unseen attacks. Ablation studies further highlight that our method can significantly improve the performance of the purifier model on robust classification. Despite the large improvements achieved, our method has a major limitation: AToP requires training on the purifier model, and as the complexity of the purifier model increases, so does the training cost. However, exploring this direction presents an intriguing research opportunity for developing a more efficient defense method by combining AT and AP.